# Parity-sensitive inhomogeneous dephasing of macroscopic spin ensembles

### v1

@@ III.1 Adding the stabilization dynamics
So far, we have only considered the free evolution governed by the dephasing Hamiltonian $H_{0}$. We now study numerically the full dissipative evolution governed by the master equation (1), using the QuTiP package [29]. Fig. 2 shows the mean fidelities $\overline{F(t)}$ for the cases of even and odd cat states, for $N=8$. In each case, we initialize the system in the even/odd cat state and measure the fidelity with respect to the initial state. In the absence of inhomogeneous broadening, the steady states achieve fidelities of $0.998$ and $0.995$ for the even and odd cat states respectively. We see that the odd cat state is significantly more susceptible to the dephasing effects. For example, at time $\Gamma_{2}t\approx 200$ with $\delta/\Gamma_{2}=10^{-2}$, the even cat state has a fidelity around $0.74$ while the fidelity of the odd cat state is only around $0.1$. This shows that the parity-sensitive dephasing derived for the simple case of only free Hamiltonian evolution applies even with the driven-dissipative stabilization terms. In Ref. [27], it was shown numerically that the even cat state |$\ket{\text{Cat}_{+}}$ is robust against inhomogeneous broadening (and other imperfections). This is consistent with our findings. We argue here that the same robustness does not apply to the odd cat state, even at larger system sizes beyond the reach of numerical simulations. For larger $N$, one expects that the effects of inhomogeneous broadening are suppressed due to stronger collective interactions, but the separation of timescales in the fidelity decay between the even and odd cat state should remain valid.
@@ IV Mean-field analysis: Synchronization phase transition
Since a direct analytical treatment of Eq. (1) is not feasible, we consider instead a mean-field approximation where we assume a product state ansatz for $\rho$. The spin coherent state in Eq. (5) is a product state, so we expect this to be a reasonable approximation for large $N$. The mean-field equations can be physically interpreted as describing the synchronization dynamics of $N$ globally coupled classical spins, containing both dissipative and coherent (also called reactive) couplings. As we will show, this system exhibits a synchronization transition with both $\eta/\Gamma_{2}$ and $\delta/\Gamma_{2}$, where the spins become desynchronized beyond a critical parameter value. This provides important insights on the robustness of the spin coherent states to inhomogeneous broadening with driven-dissipative stabilization, for large system sizes.
The mean-field equations governing the dynamics of the $m$-th spin read (see Appendix A for a detailed derivation):
@@
Figure 4: Steady state value of the phase spread $\zeta$ against the dimensionless detuning $\tilde{\delta}=N^{2}\delta/\Gamma_{2}$ and dimensionless squeezing $\tilde{\eta}=\eta/\Gamma_{2}$, for $N=10^{4}$ spins. The white region denote data points where the spins fail to converge to a synchronized state at long times. The red line marks the elliptical boundary of the synchronization region given by Eq. (22), with fitting parameters $a\approx 0.03125$ and $b\approx 2.2077$. For small values of $\tilde{\eta}$, some points within the synchronization region fail to converge to the steady state within the simulation time due to the slow dynamics.
@@ IV.2 Oppositely detuned sub-ensembles
$\begin{split}\frac{\dot{A}}{A}&=-2\eta z\left(\cos^{2}\zeta-\frac{1}{N}\cos 2\zeta\right)\\ &+\Gamma_{2}z\left(\cos^{2}\zeta-\frac{1}{N}\right)\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\\ \dot{\zeta}&=\delta+\eta z\left(1-\frac{2}{N}\right)\sin 2\zeta\\ &-\frac{\Gamma_{2}}{2}z\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\sin 2\zeta\\ \frac{\dot{z}}{A^{2}}&=8\eta\left(\cos^{2}\zeta-\frac{1}{N}\cos 2\zeta\right)\\ &-4\Gamma_{2}\left(\cos^{2}\zeta-\frac{1}{N}\right)\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\\ &-4\Gamma_{2}(1+z)\cos^{2}\zeta\end{split}$ (20)
To verify the accuracy of Eq. (20), we compare numerical simulations of Eq. (20) with the full mean-field equations (16) and (17). It is also simple to see that this is consistent with Eq. (19) by setting $\delta=0,\phi\approx\pi/4,z\approx-1$ and taking the large $N$ limit.
Compared to the previous case of identical frequencies, this set of equations is much harder to solve analytically. For small $\zeta$, we can get an approximate solution by expanding Eq. (20) to first order in $\zeta$ and also first order in $1/N$, the approximate steady state solution corresponding to the synchronized state is (see Appendix B)
@@
where $\tilde{\delta}\equiv N^{2}\delta/\Gamma_{2}$ and $\tilde{\eta}\equiv\eta/\Gamma_{2}$ are the dimensionless frequency detuning and squeezing respectively, and $a,b$ are the fitting parameters. The threshold values are thus $\tilde{\eta}_{c}=2a$ and $\tilde{\delta}_{c}=ab$. From numerical fitting, we find that $a\approx 0.03125$ and $b\approx 2.2077$ which gives $\tilde{\eta}_{c}\approx 0.0625N$ and $\tilde{\delta}_{c}\approx 0.06888N^{2}$. The value of $\tilde{\eta}_{c}$ agrees excellently with the theoretical prediction of $N/16$.
Since $\tilde{\delta}_{c}\propto N^{2}$, the critical detuning $\delta_{c}/\Gamma_{2}\approx 0.06888$ is independent of $N$. This suggests that the robustness of the spin coherent state to inhomogeneous broadening is independent of $N$, for a fixed $\eta/\Gamma_{2}$. Since the odd cat state is more fragile than the spin coherent state by the parity-sensitive dephasing effect, this supports the lack of robustness for the odd cat state, consistent with the numerical results in Fig. 2 for the quantum model.
@@ V Discussion and Outlook
In this work, we analyze the effects of symmetry-breaking inhomogeneous dephasing on collective spin states such as spin coherent states and cat states $|\text{Cat}_{\pm}\rangle$ formed from the superposition of two spin coherent states offset by a $\pi$-phase. We show that the spin dephasing is highly sensitive to the parity symmetry of the collective spin state: For $|\text{Cat}_{+}\rangle$, even symmetry provides enhanced protection against inhomogeneous dephasing which grows with the system size $N$. On the contrary, the odd symmetry in $|\text{Cat}_{-}\rangle$ renders it significantly more fragile to inhomogeneous dephasing, causing the fidelity to decay rapidly.
This has important implications in using macroscopic spin ensembles as ‘cat qubits’ to encode logical quantum information, motivated by recent successful experiments in superconducting quantum circuits. Due to the collectively enhanced interactions in spin ensembles scaling as $\sqrt{N}$, they stand to benefit from greater error-correcting capabilities compared to their bosonic counterpart. However, using spin ensembles also come with a different set of experimental challenges like particle loss which has to be carefully addressed. In certain implementations such as rare-earth ions, inhomogeneous broadening arising from spatially-varying magnetic fields break permutation symmetry which is required for the spins to behave collectively, and causes spin dephasing. In Ref. [27], it was demonstrated explicitly that $|\text{Cat}_{+}\rangle$ is robust against such inhomogeneous dephasing, which is consistent with our results. However, our results also suggest that the odd cat state $|\text{Cat}_{-}\rangle$ is significantly more fragile. This means that the lifetime of $|\text{Cat}_{-}\rangle$ is limited by the dephasing timescale from the inhomogeneous broadening, which is on the order of $1$ $\mu\text{s}$ (taking $\delta\sim 1$ MHz). Consequently, encoding the logical qubit using $|\text{Cat}_{\pm}\rangle$ is unlikely to result in logical lifetimes significantly longer than the break-even point.
@@ Acknowledgments
L.C.K. acknowledges support from the Ministry of Education, Singapore and the National Research Foundation, Singapore. S.T acknowledges support from the National Research Foundation, Singapore. The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.
Appendix A Derivation of the semiclassical mean-field equations
@@
Appendix B Two detuned ensembles, small $\delta$ limit
When the detuning $\delta$ is small, we expect the steady state value for the phase $\phi$ to be close to $\pi/4$, as was shown in Eq. (19). Writing $\phi=\pi/4\pm\zeta$, we obtain the mean-field equations (20). Expanding the equations to linear order in $\zeta$, and then to order $1/N$, we obtain the simplifed equations (setting $\Gamma_{2}=1$ for notational simplicity)

### v2

@@ III.2 Adding the stabilization dynamics
So far, we have only considered the free evolution governed by the dephasing Hamiltonian $H_{0}$. We now study numerically the full dissipative evolution governed by the master equation (1), using the QuTiP package [30]. Fig. 2 shows the mean fidelities $\overline{F(t)}$ for the cases of even and odd cat states, for $N=8$. In each case, we initialize the system in the even/odd cat state and measure the fidelity with respect to the initial state. In the absence of inhomogeneous broadening, the steady states achieve fidelities of $0.998$ and $0.995$ for the even and odd cat states respectively. We see that the odd cat state is significantly more susceptible to the dephasing effects. For example, at time $\Gamma_{2}t\approx 200$ with $\delta/\Gamma_{2}=10^{-2}$, the even cat state has a fidelity around $0.74$ while the fidelity of the odd cat state is only around $0.1$. This shows that the parity-sensitive dephasing derived for the simple case of only free Hamiltonian evolution applies even with the driven-dissipative stabilization terms. In Ref. [28], it was shown numerically that the even cat state |$\ket{\text{Cat}_{+}}$ is robust against inhomogeneous broadening (and other imperfections). This is consistent with our findings. We argue here that the same robustness does not apply to the odd cat state, even at larger system sizes beyond the reach of numerical simulations. For larger $N$, one expects that the effects of inhomogeneous broadening are suppressed due to stronger collective interactions, but the separation of timescales in the fidelity decay between the even and odd cat state should remain valid, at small amplitudes.
@@ IV Mean-field analysis: Synchronization phase transition
Since a full analytical treatment of Eq. (1) is not feasible, we consider instead a mean-field approximation where we assume a product state ansatz for $\rho$. The spin coherent state in Eq. (5) is a product state, so we expect this to be a reasonable approximation for large $N$. However, note that the mean-field approximation cannot capture superpositions and thus does not account for parity-sensitive dephasing. Nonetheless, for large amplitudes, the discrepancy in the dephasing effects between |$\ket{\text{Cat}_{\pm}}$ becomes irrelevant, and behaves similarly as the spin coherent state. Thus, we expect the study of the mean-field model to yield useful insights about the stability of the spin cat states in the full quantum model (1).
The mean-field equations can be physically interpreted as describing the synchronization dynamics of $N$ globally coupled classical spins, containing both dissipative and coherent (also called reactive) couplings. As we will show, this system exhibits a synchronization transition with both $\eta/\Gamma_{2}$ and $\delta/\Gamma_{2}$, where the spins become desynchronized beyond a critical parameter value. This provides important insights on the robustness of the spin coherent states to inhomogeneous broadening with driven-dissipative stabilization, for large system sizes.
The mean-field equations governing the dynamics of the $m$-th spin read (see Appendix B for a detailed derivation):
@@ IV.2 Oppositely detuned sub-ensembles
$\begin{split}\frac{\dot{A}}{A}&=-2\eta z\left(\cos^{2}\zeta-\frac{1}{N}\cos 2\zeta\right)\\ &+\Gamma_{2}z\left(\cos^{2}\zeta-\frac{1}{N}\right)\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\\ &-2\Gamma_{2}A^{2}\cos^{2}\zeta,\end{split}$ (26)
$\begin{split}\dot{\zeta}&=\delta+\eta z\left(1-\frac{2}{N}\right)\sin 2\zeta\\ &-\frac{\Gamma_{2}}{2}z\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\sin 2\zeta,\end{split}$ (27)
and
$\begin{split}\frac{\dot{z}}{A^{2}}&=8\eta\left(\cos^{2}\zeta-\frac{1}{N}\cos 2\zeta\right)\\ &-4\Gamma_{2}\left(\cos^{2}\zeta-\frac{1}{N}\right)\left[1+z+NA^{2}\left(1-\frac{2}{N}\right)\cos^{2}\zeta\right]\\ &-4\Gamma_{2}(1+z)\cos^{2}\zeta.\end{split}$ (28)
To verify the accuracy of Eqs. (26), (27), and (28), we numerically simulate them and compare the results with the full mean-field equations (22) and (23). It is also simple to see that this is consistent with Eq. (25) by setting $\delta=0,\phi\approx\pi/4,z\approx-1$ and taking the large $N$ limit.
Compared to the previous case of identical frequencies, this set of equations is much harder to solve analytically. For small $\zeta$, we can get an approximate solution by expanding Eqs. (26), (27), and (28) to first order in $\zeta$ and also first order in $1/N$. The approximate steady state solution corresponding to the synchronized state is (see Appendix C)
@@
Figure 4: Steady state value of the phase spread $\zeta$ against the dimensionless detuning $\tilde{\delta}=N^{2}\delta/\Gamma_{2}$ and dimensionless squeezing $\tilde{\eta}=\eta/\Gamma_{2}$, for $N=10^{4}$ spins. The white region denote data points where the spins fail to converge to a synchronized state at long times. The red line marks the elliptical boundary of the synchronization region given by Eq. (30), with fitting parameters $a\approx 0.03125N$ and $b\approx 2.2077N$. For small values of $\tilde{\eta}$, some points within the synchronization region fail to converge to the steady state within the simulation time due to the slow dynamics.
@@
where $\tilde{\delta}\equiv N^{2}\delta/\Gamma_{2}$ and $\tilde{\eta}\equiv\eta/\Gamma_{2}$ are the dimensionless frequency detuning and squeezing respectively, and $a,b$ are the fitting parameters. The threshold values are thus $\tilde{\eta}_{c}=2a$ and $\tilde{\delta}_{c}=ab$. From numerical fitting, we find that $a\approx 0.03125N$ and $b\approx 2.2077N$ which gives $\tilde{\eta}_{c}\approx 0.0625N$ and $\tilde{\delta}_{c}\approx 0.06888N^{2}$. The value of $\tilde{\eta}_{c}$ agrees excellently with the theoretical prediction of $N/16$.
The mean-field analysis suggests that the spin coherent state is robust against inhomogeneous broadening, under the condition that the broadening linewidth does not scale with $N$. From Eq. (29), we have $\zeta\approx N^{2}\delta/4|\alpha|^{4}\Gamma_{2}$. This implies that $\delta/\Gamma_{2}$ needs to scale as $\sim 1/N^{2}$ for stability at a constant amplitude $|\alpha|$. This condition is satisfied for constant $\delta$ since $\Gamma_{2}\sim N^{2}$ (see Appendix A). A full quantum analysis of the robustness remains an open problem and is left as future work.
@@ V Discussion and Outlook
In this work, we analyze the effects of symmetry-breaking inhomogeneous dephasing on collective spin states such as spin coherent states and cat states $|\text{Cat}_{\pm}\rangle$ formed from the superposition of two spin coherent states offset by a $\pi$-phase.
To summarize the core arguments of our paper:
  1. 1.
    For cat states with small amplitudes $(|\alpha|\ll 1)$, the even spin cat state is more robust to inhomogeneous dephasing than the spin coherent state of the same amplitude, which is in turn more robust than the odd spin cat state. We demonstrate this parity-sensitive dephasing analytically without dissipative stabilization, which is supported with exact numerics of the Lindblad master equation for small system sizes.
  2. 2.
    However, for large cat states $|\alpha|\gg 1$, the sensitivity to the number parity vanishes. Consequently, the robustness of the spin cat states is similar to that of a spin coherent state with the same $|\alpha|$. Thus, we can estimate the stability of large spin cat states by analyzing the spin coherent state.
  3. 3.
    The stability of the spin coherent state (with dissipative stabilization) can be analyzed using a mean-field approach to the Lindblad master equation. The quantum model is now approximated by a synchronization model of classical spins.
  4. 4.
    We then show analytically that the mean-field model exhibits a synchronization phase transition at a critical inhomogeneous broadening linewidth, beyond which the synchronization of the classical spins breaks down. We argue that this characterizes the robustness of the spin coherent states (and from point 2, the large cat states of any parity as well).
  5. 5.
    By analyzing the robustness of the classical synchronization model, we estimate that the dissipatively stabilized spin cats (with amplitude $|\alpha|\gg 1$) are robust to inhomogeneous broadening if $\delta/\Gamma_{2}\sim 1$. This is satisfied if the broadening linewidth $\delta$ does not increase proportionally to $N$, since $\Gamma_{2}\sim N$ in a realistic implementation of the model.
Our findings have important implications in using macroscopic spin ensembles as ‘cat qubits’ to encode logical quantum information, motivated by recent successful experiments in superconducting quantum circuits. Due to the collectively enhanced interactions in spin ensembles scaling as $\sqrt{N}$, they stand to benefit from greater error-correcting capabilities compared to their bosonic counterpart. However, using spin ensembles also come with a different set of experimental challenges like particle loss which has to be carefully addressed. In certain implementations such as rare-earth ions, inhomogeneous broadening arising from spatially-varying magnetic fields break permutation symmetry which is required for the spins to behave collectively, and causes spin dephasing. In Ref. [28], it was demonstrated explicitly that $|\text{Cat}_{+}\rangle$ is robust against such inhomogeneous dephasing, which is consistent with our results. However, our results also suggest that the odd cat state $|\text{Cat}_{-}\rangle$ is significantly more fragile. This means that the lifetime of $|\text{Cat}_{-}\rangle$ is limited by the dephasing timescale from the inhomogeneous broadening, which is on the order of $1$ $\mu\text{s}$ (taking $\delta\sim 1$ MHz). Consequently, encoding the logical qubit using $|\text{Cat}_{\pm}\rangle$ is unlikely to result in logical lifetimes significantly longer than the break-even point, unless the bosonic amplitude is large, in which our results indicate that |$\ket{\text{Cat}_{-}}$ can be potentially robust.
@@ Acknowledgments
We thank the anonymous referees for helpful suggestions. L.C.K. acknowledges support from the Ministry of Education, Singapore and the National Research Foundation, Singapore. S.T acknowledges support from the National Research Foundation, Singapore. The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.
Appendix A Derivation of the spin master equation
Here, we provide a physical derivation of the master equation (1) in the main text. Our approach provides an alternative to that in Ref. [28], based on time-averaging of the Hamiltonian.
Consider the Hamiltonian
$H=\sum_{n=1}^{N}\frac{\omega_{0}+\delta_{n}}{2}\sigma_{n}^{z}+2\omega_{0}a^{\dagger}a+g\sum_{n=1}^{N}(a^{\dagger}\sigma_{n}^{-}+a\sigma_{n}^{+})+\chi^{(2)}(a+a^{\dagger})^{3},$ (31)
which describes the coupling between the spins and the cavity mode with bosonic annihilation operator $a$, in the presence of a nonlinear medium with a large second-order susceptibility $\chi^{(2)}$. The cavity is assumed to be strongly dissipative, i.e., low quality factor. In the dispersive regime where the cavity-spin detuning $\Delta\approx\omega_{0}$ is large, i.e., $g/\Delta\ll 1$, we have $a\to a+(g/\Delta)\sum_{n}\sigma_{n}^{-}+O(g^{2}/\Delta^{2})$, and the Hamiltonian can be approximated by (in the rotating frame at frequencies $2\omega_{0}$ with respect to the cavity and $\omega_{0}$ with respect to the spins)
$H\approx\sum_{n=1}^{N}\frac{\delta_{n}}{2}\sigma_{n}^{z}+g_{2}\sum_{m,n=1}^{N}(a\sigma_{m}^{-}\sigma_{n}^{-}+a^{\dagger}\sigma_{m}^{+}\sigma_{n}^{+})$ (32)
where
$g_{2}=\chi^{(2)}\left(\frac{g}{\Delta}\right)^{2}.$ (33)
Note that $\chi^{(2)}$ must be sufficiently large such that the omitted terms are negligible. Next, we go into the displaced frame $a\to a+\alpha_{d}$, which can be realized by applying a classical drive with amplitude $\alpha_{d}$ to the cavity. This gives
$H=\sum_{n=1}^{N}\frac{\delta_{n}}{2}\sigma_{n}^{z}+g_{2}\sum_{m,n=1}^{N}(\alpha_{d}\sigma_{m}^{-}\sigma_{n}^{-}+\alpha_{d}^{*}\sigma_{m}^{+}\sigma_{n}^{+})+g_{2}\sum_{m,n=1}^{N}(a\sigma_{m}^{-}\sigma_{n}^{-}+a^{\dagger}\sigma_{m}^{+}\sigma_{n}^{+}).$ (34)
In the regime where the dissipation rate of the cavity $\kappa$ is much larger than $g_{2}$, the cavity mode can be adiabatically eliminated. This results in an effective master equation describing the driven-dissipative dynamics of the spins mediated by the cavity:
$\dot{\rho}=-i[H_{\text{spin}},\rho]+\frac{4g_{2}^{2}}{\kappa}\mathcal{D}[\sum_{m,n}\sigma_{m}^{-}\sigma_{n}^{-}]\rho$ (35)
with the spin Hamiltonian
$H_{\text{spin}}=\frac{1}{2}\sum_{n=1}^{N}\delta_{n}\sigma_{n}^{z}+g_{2}\sum_{m,n=1}^{N}(\alpha_{d}\sigma_{m}^{-}\sigma_{n}^{-}+\alpha_{d}^{*}\sigma_{m}^{+}\sigma_{n}^{+}).$ (36)
This corresponds to the master equation (1) in the main text, with parameters
$\frac{\eta e^{i\varphi}}{N}=g_{2}\alpha_{d},$ (37)
and
$\frac{\Gamma_{2}}{N^{2}}=\frac{4g_{2}^{2}}{\kappa}.$ (38)
The cat amplitude is
$|\alpha|^{2}=\frac{2\eta}{\Gamma_{2}}=\frac{\alpha_{d}\kappa}{2g_{2}N}.$ (39)
This implies that the classical drive amplitude $\alpha_{d}\sim N$ should scale proportionally to the number of spins, in order to keep the cat amplitude fixed.
Appendix B Derivation of the semiclassical mean-field equations
@@
Appendix C Two detuned ensembles, small $\delta$ limit
When the detuning $\delta$ is small, we expect the steady state value for the phase $\phi$ to be close to $\pi/4$, as was shown in Eq. (25). Writing $\phi=\pi/4\pm\zeta$, we obtain the mean-field equations (26), (27), and (28). Expanding the equations to linear order in $\zeta$, and then to order $1/N$, we obtain the simplifed equations (setting $\Gamma_{2}=1$ for notational simplicity)